2. To go to a nearby convenience store from her dormitory, a student walks 12.0 m, north and 5.0 m, east. What is the student’s displacement from her dormitory? (NB. Displacement is a vector quantity hence you need to determine its specific direction)
1
Expert's answer
2021-10-25T17:55:14-0400

The magnitude is


"d=\sqrt{12^2+5^2}=13\text{ m},\\\space\\
\theta=\arctan\frac{12}{5}=67°\text{ N of E}."